40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 2 16:29:40 2020, gautam, Update, Electronics, c1psl VME crate removed, Acro-crate installed c1psl.pdf
    Reply  Tue Mar 3 17:20:14 2020, gautam, Update, Electronics, More cabling removed 
       Reply  Wed Mar 4 16:18:31 2020, gautam, Update, Electronics, More cabling removed 1X2EuroBefore.JPGIOO.png
          Reply  Wed Mar 4 21:02:49 2020, Koji, Update, Electronics, More cabling removed 
    Reply  Tue Mar 3 17:59:33 2020, Yehonathan, Update, Electronics, PSL Shutter and PMC TRANSPD working 
       Reply  Thu Mar 5 15:03:48 2020, Yehonathan, Update, Electronics, PSL Shutter and PMC TRANSPD working 
          Reply  Thu Mar 5 19:45:23 2020, Jon, Summary, PSL, C1PSL in-situ test results c1psl_feedthrough_wiring_-_By_Connector_(3).pdf
             Reply  Wed Mar 11 18:12:53 2020, gautam, Summary, PSL, WFS Demod board modifications D980233-B_Mar2020Mods.pdf
Message ID: 15243     Entry time: Tue Mar 3 17:59:33 2020     In reply to: 15238     Reply to this: 15255
Author: Yehonathan 
Type: Update 
Category: Electronics 
Subject: PSL Shutter and PMC TRANSPD working 

I used existing BNC cables running from the PSL table to the PSL rack and reassigned them to the PSL Shutter and PMC transmission PD channels.

The PSL shutter turned out to be a sinking channel. Jordan reconnected the PSL shutter wires to a sinking BIO Acromag. Channel list is updated.

Both channels have been tested to be working as expected.


gautam add on about EPICS:

  • the PSL shutter channels were previously hosted on c1aux.
  • I didn't comment out the original database entries on c1aux because we changed the prefix for all these channels - i.e. C1:AUX-PSL_Shutter --> C1:PSL-PSL_Shutter.
  • Modified the LSC offset script to close/open the PSL shutter by writing to the correct channel now.
  • there is some EPICS logic that checks the main volume pressure and prevents the opening of the PSL shutter if the main volume pressure is between 0.003 torr and 500 torr. I preserved this capability (so there are some associated soft channels in the database as well).

P.S - there is a problem we noticed - if the modbus process is started with the local subnet not having a fixed IP address, then all the EPICS channels will not be responsive. The way to fix this is to run the following sequence of commands:

sudo systemctl stop modbusIOC.service
sudo ifdown enp4s0
sudo ifup enp4s1
sudo ssytemctl start modbusIOC.service
ELOG V3.1.3-